# SCOPE: Stochastic Cartographic Occupancy Prediction Engine for Uncertainty-Aware Dynamic Navigation

### v1

@@ VI-C Simulation Results
Table III summarizes these navigation results, where we observe three key phenomena. First, compared to DWA series variants (i.e., DWA, DWA-DeepTracking-P, DWA-SCOPE-P, and DWA-SO-SCOPE-P), our proposed DWA-SCOPE-PU and DWA-SO-SCOPE-PU policies have the highest success rate in each crowd size, while DWA-SCOPE-PU has almost the shortest path length. This shows that the prediction costmap from our OGM predictors is able to help the currently existing model-based navigation policy (i.e., DWA) to provide safer and shorter paths, and combining it with its associated uncertainty costmap can achieve a much better navigation performance (i.e., PU is better than P alone). Furthermore, it shows that our software-optimized SO-SCOPE predictor can maintain good robot navigation performance even though it leads to longer paths compared to the SCOPE predictor. All these results demonstrates the effectiveness of our proposed SCOPE series predictors in helping design safe robot navigation policies in crowded dynamic scenes.
The reasons why the OGM prediction and its uncertainty information can improve robot safe navigation can be explained qualitatively through Fig. 16. It shows the difference of nominal paths and costmaps generated by four different planners (i.e., DWA, DWA-SCOPE-P, DWA-SCOPE-PU, and DWA-SO-SCOPE-PU) in the simulated lobby environment. The default DWA planner [8] only cares about the current state of the environment and generates a costmap based on the perceived obstacles. The predictive DWA planner (i.e., DWA-SCOPE-P) using the prediction map of our proposed SCOPE predictor can generate a costmap with predicted obstacles. The predictive uncertainty-aware DWA planner (i.e., DWA-SCOPE-PU and DWA-SO-SCOPE-PU) using both the prediction map and uncertainty map of our proposed SCOPE predictor can generate a safer costmap with predicted obstacles and uncertainty regions. Note that the path generated by DWA-SO-SCOPE-PU is more tortuous than that of DWA-SCOPE-PU, which may explain why DWA-SO-SCOPE-PU requires longer paths to achieve a similar success rate as DWA-SCOPE-PU. These additional predicted obstacles and uncertainty regions of our proposed DWA-SCOPE-PU and DWA-SO-SCOPE-PU planners enable the robot to follow safer nominal paths and reduce collisions with obstacles, especially moving pedestrians. See the accompanying Multimedia for a detailed simulation navigation demonstration.
@@
(b) DWA-SCOPE-P
(c) DWA-SCOPE-PU
(d) DWA-SO-SCOPE-PU
Figure 16: Robot reactions and their corresponding costmaps generated by different DWA-based control policies in the simulated lobby environment. The robot (black disk) is avoiding pedestrians (colorful square boxes) and reaching the goal (red disk) according to the nominal path (green line) planned by the costmap (square white map).
Second, our proposed DWA-SCOPE-P and DWA-SO-SCOPE-P have a higher success rate than DWA-DeepTracking-P at every crowd size, indicating that higher prediction accuracy can also improve robot safe navigation performance. Third, compared with all other state-of-the-art control policies, our proposed DRL-VO-SO-SCOPE-PU policy has the highest success rate and the fastest average speed in every situation. This suggests that our software-optimized SO-SCOPE predictor using a simple knowledge distillation network and prediction uncertainty statistical lookup table can also provide additional benefits for the learning-based robot safe navigation performance in crowded dynamic environments with varying crowd densities. Furthermore, it shows the potential capability of our software-optimized predictive uncertainty-aware navigation framework for different high computational load learning-based control policies.
@@
Figure 17: Robot deployed with DWA-SCOPE-PU reactions to moving pedestrians in the indoor hallway with the high crowd density at different times.
@@
Figure 18: Robot deployed with DRL-VO-SO-SCOPE-PU reactions to moving pedestrians in the indoor hallway with the high crowd density at different times.
@@ VI-D Hardware Results
Besides the simulated experiments, we also conduct a real-world experiment to demonstrate the applicability of our DWA-SCOPE-PU and DRL-VO-SO-SCOPE-PU policies in the real world. Both control policies use the amcl ROS package to provide the robot localization service in known maps. Note that our DRL-VO-SO-SCOPE-PU control policy requires additional running of three deep learning networks (i.e., YOLOv3 [79], SO-SCOPE, and DRL-VO) and one optimization-based multiple hypothesis tracker (MHT) [80] to implement corresponding pedestrian detection, environment OGM prediction, navigation control, and pedestrian tracking.
From the attached Multimedia and Fig. 17, we can see how our robot deployed with DWA-SCOPE-PU is able to actively avoid collisions with walking students crossing the hallway, safely avoid static students standing in the hallway, and reach predefined goals by following predictive uncertainty-ware nominal paths, traveling a total length of 76.10 m and an average speed of 0.42 m/s. It demonstrates the real-world effectiveness of our proposed SCOPE predictors and DWA-SCOPE-PU planner.
In addition, from the attached Multimedia and Fig. 18, we can see that even with three learning-based blocks, our robot deployed with DRL-VO-SO-SCOPE-PU is still able to quickly and actively avoid collisions with walking students crossing the hallway and reach predefined goals by following predictive uncertainty-ware nominal paths, traveling a total length of 86.41 m and an average speed of 0.47 m/s. It demonstrate our software-optimized SO-SCOPE predictor is hardware friendly and our software-optimized predictive uncertainty-aware navigation framework can be combined with different high computational load learning-based algorithms.
@@ VII Conclusion
In this article, we propose a family of hardware-friendly stochastic occupancy grid map prediction algorithms (i.e., SCOPE++, SCOPE, and SO-SCOPE) that provides mobile robots with the ability to accurately and robustly predict the future states of complex, dynamic, human-occupied environments. Specifically, we first propose two VAE-based stochastic predictors (i.e., SCOPE++ and SCOPE) that exploit information from robot motion, object motion, and static scene geometry, and use a VAE-based network to fuse these useful information into the latent space and predict the distribution of future environment states. We then perform software optimization on them using knowledge distillation and uncertainty quantification operations to propose a hardware-friendly SO-SCOPE to extend their practical applications in resource-constrained robots. Furthermore, we propose a novel predictive uncertainty-aware navigation framework by combining these proposed stochastic predictors with the costmap-based ROS navigation stack to improve the performance of current state-of-the-art model-based and learning-based control policies for safe navigation in crowded dynamic scenes.
We demonstrate that our proposed SCOPE family achieves smaller absolute error, higher structure similarity, and higher tracking accuracy than the other state-of-the-art image-based predictors on three different simulated and real-world datasets collected by three different robot models. It also provides a range of plausible and diverse future prediction states for complex stochastic environments. We further demonstrate through operational analysis and experiments on embedded computing devices that our proposed software-optimized SO-SCOPE prediction engine is faster and more memory-efficient, allowing it to provide real-time inference in resource-constrained robots. Lastly, we demonstrate through simulated and hardware experiments that we can easily integrate the predicted maps into the existing navigation framework, all of which benefit from leveraging the prediction and uncertainty information of our SCOPE family. In summary, our SCOPE predictors are hardware-friendly and improve mobile robots’ ability to safely navigate through crowded dynamic environments.

### v2

@@ VI-C Simulation Results
Table III summarizes these navigation results, where we observe three key phenomena. First, compared to DWA series variants (i.e., DWA, DWA/DeepTracking/P, DWA/SCOPE/P, and DWA/SO-SCOPE/P), our proposed DWA/SCOPE/PU and DWA/SO-SCOPE/PU policies have the highest success rate in each crowd size, while DWA/SCOPE/PU has almost the shortest path length. This shows that the prediction costmap from our OGM predictors is able to help the currently existing model-based navigation policy (i.e., DWA) to provide safer and shorter paths, and combining it with its associated uncertainty costmap can achieve a much better navigation performance (i.e., PU is better than P alone). Furthermore, it shows that our software-optimized SO-SCOPE predictor can maintain good robot navigation performance even though it leads to longer paths compared to the SCOPE predictor. All these results demonstrate that our proposed SCOPE series predictors can improve safe robot navigation in crowded dynamic scenes.
The reasons why the OGM prediction and its uncertainty information can improve robot safe navigation can be explained qualitatively through Fig. 13. It shows the difference of nominal paths and costmaps generated by four planners (i.e., DWA, DWA/SCOPE/P, DWA/SCOPE/PU, and DWA/SO-SCOPE/PU) in the simulated lobby environment. The default DWA planner [10] only cares about the current state of the environment and generates a costmap based on the perceived obstacles. The predictive DWA planner (i.e., DWA/SCOPE/P) using the prediction map of our proposed SCOPE predictor can generate a costmap with predicted obstacles. The predictive uncertainty-aware DWA planner (i.e., DWA/SCOPE/PU and DWA/SO-SCOPE/PU) using both the prediction map and uncertainty map of our proposed SCOPE predictor can generate a safer costmap with predicted obstacles and uncertainty regions. Note that the path generated by DWA/SO-SCOPE/PU is more tortuous than that of DWA/SCOPE/PU, which may explain why DWA/SO-SCOPE/PU requires longer paths to achieve a similar success rate as DWA/SCOPE/PU. These additional predicted obstacles and uncertainty regions of our proposed DWA/SCOPE/PU and DWA/SO-SCOPE/PU planners enable the robot to follow safer nominal paths and reduce collisions with obstacles, especially moving pedestrians. See the accompanying Multimedia for a detailed simulation navigation demonstration.
@@
(b) DWA/SCOPE/P
(c) DWA/SCOPE/PU
(d) DWA/SO-SCOPE/PU
Figure 13: Robot reactions and their final costmaps generated by different DWA-based control policies in the simulated lobby environment, which shows that the final master costmap generated by the prediction map and its uncertainty map can provide safer path planning. The robot (black disk) avoids pedestrians (colorful square boxes, each color represents a pedestrian) and reaches the goal (red disk) according to the nominal path (green curve) and local path (red curve) planned by the costmap (square grey map).
Second, our proposed DWA/SCOPE/P and DWA/SO-SCOPE/P have a higher success rate than DWA/DeepTracking/P at every crowd size, indicating that higher prediction accuracy can also improve safe navigation performance. Third, compared with all other state-of-the-art control policies, our proposed DRL-VO/SO-SCOPE/PU policy has the highest success rate and the fastest average speed in every situation. This suggests that our software-optimized SO-SCOPE predictor using a simple knowledge distillation network and prediction uncertainty statistical lookup table can also provide additional benefits for the learning-based safe navigation performance in crowded dynamic environments with varying crowd densities. Furthermore, it shows the potential capability of our software-optimized predictive uncertainty-aware navigation framework for different high-computational load learning-based policies.
@@
Figure 14: Robot deployed with DWA/SCOPE/PU reactions to moving pedestrians in the indoor hallway with the high crowd density at different times.
@@
Figure 15: Robot deployed with DRL-VO/SO-SCOPE/PU reactions to moving pedestrians in the indoor hallway with the high crowd density at different times.
@@ VI-D Hardware Results
Besides simulated experiments, we also conduct real-world experiments to demonstrate the applicability of our DWA/SCOPE/PU and DRL-VO/SO-SCOPE/PU policies. Both control policies use the amcl ROS package to provide the robot localization in known maps. Note that our DRL-VO/SO-SCOPE/PU control policy requires additional running of three deep learning networks (i.e., YOLOv3 [78], SO-SCOPE, and DRL-VO [13]) and one optimization-based multiple hypothesis tracker (MHT) [79] to implement corresponding pedestrian detection, environment OGM prediction, navigation control, and pedestrian tracking. From the attached Multimedia and Fig. 14, we can see how our robot deployed with DWA/SCOPE/PU can actively avoid collisions with walking students crossing the hallway, safely avoid static students standing, and reach predefined goals by following predictive uncertainty-aware nominal paths, traveling a total length of 76.10 m and an average speed of 0.42 m/s. It demonstrates the real-world effectiveness of our proposed SCOPE predictor and DWA/SCOPE/PU planner.
In addition, from the attached Multimedia and Fig. 15, we can see that even with three learning-based blocks, our robot deployed with DRL-VO/SO-SCOPE/PU is still able to quickly and actively avoid collisions with walking students crossing the hallway and reach predefined goals by following predictive uncertainty-ware nominal paths, traveling a total length of 86.41 m and an average speed of 0.47 m/s. It demonstrates our software-optimized SO-SCOPE predictor is hardware friendly and our software-optimized predictive uncertainty-aware navigation framework can be combined with different high computational load learning-based algorithms.
@@ VII Conclusion
In this article, we propose a family of hardware-friendly stochastic occupancy grid map prediction algorithms (i.e., SCOPE++, SCOPE, and SO-SCOPE) that provides mobile robots with the ability to accurately and robustly predict the future states of complex, dynamic, human-occupied environments. Specifically, we first propose two VAE-based stochastic predictors (i.e., SCOPE++ and SCOPE) that exploit information from robot motion, object motion, and static scene geometry, and use a VAE-based network to fuse this useful information into the latent space and predict the distribution of future environment states. We then perform software optimization on them using knowledge distillation and uncertainty quantification operations to propose a hardware-friendly SO-SCOPE, which significantly improves the inference speed, addresses their sampling-based memory-intensive nature, and extends their practical applications in resource-constrained robots. Furthermore, we propose a novel predictive uncertainty-aware navigation framework by combining these proposed stochastic predictors with the costmap-based ROS navigation stack to improve the performance of current state-of-the-art model-based and learning-based control policies for safe navigation in crowded dynamic scenes.
We demonstrate that our proposed SCOPE family achieves smaller absolute error, higher structure similarity, and higher tracking accuracy than the other state-of-the-art image-based predictors on three different simulated and real-world datasets collected by three different robot models. It also provides a range of plausible and diverse future prediction states for complex stochastic environments. We further demonstrate through operational analysis and experiments on embedded computing devices that our proposed software-optimized SO-SCOPE prediction engine is much faster and more memory-efficient, allowing it to provide real-time inference with other resource-intensive algorithms in resource-constrained robots. Lastly, we demonstrate through simulated and hardware experiments that we can easily integrate the predicted maps into the existing navigation framework, all of which benefit from leveraging the prediction and uncertainty information of our SCOPE family. In summary, our proposed SCOPE predictors are hardware-friendly and improve mobile robots’ ability to safely navigate through crowded dynamic environments.